OpenStudy (anonymous):

My Practice Space... see if this works \(\large \frac {\sqrt{3}}{2} \]

OpenStudy (anonymous):

guess not.. this is a test for inline latex \(\large \frac {\sqrt{3^5}}{2y^7} \)

OpenStudy (anonymous):

this is the quadratic formula: \(\huge x=\frac {-b \pm \sqrt{b^2-4(a)(c)}}{2a} \)

OpenStudy (anonymous):

cut and paste.... \[\int_{a}^{b}lnxdx=xlnx+C \]

OpenStudy (anonymous):

you have to open and close with the same tag so \(\large \frac {\sqrt{3}}{2} \] does not work because you open with ( but closed with ]

OpenStudy (anonymous):

\(\large \frac {\sqrt{3}}{2} \) does it in line

OpenStudy (theviper):

\[\Huge{\color{red}{\sqrt{3\over2}}}\]

OpenStudy (anonymous):

this is where you can get practice on all your latexing...

OpenStudy (anonymous):

But how can I??

OpenStudy (anonymous):

as you can see everyone post up a question and basically work/practice their latexing here. to get hints on cancelling and do any kind of things, you can click on their posts to see exactly what they did...

OpenStudy (anonymous):

Ok thanks.. I had not any clue about LATEXING...

OpenStudy (anonymous):

best thing to do is right click ---> show math as ---> tex commands to see the latex codes the used... copy/paste is usually a good start...

OpenStudy (anonymous):

\[\\Huge{\color{white}{{\heartsuit}}} \]

OpenStudy (anonymous):

\(\Huge{\color{blue}{{\heartsuit}}} \)

OpenStudy (anonymous):

\Huge{\color{blue}{{\heartsuit}}}

OpenStudy (anonymous):

\[\huge \frac{sinx}{x}=6 \]

OpenStudy (anonymous):

Why I did get that Heart??

OpenStudy (anonymous):

"\Huge{\color{blue}{{\heartsuit}}}"

OpenStudy (anonymous):

\[\Huge{\color{blue}{{\heartsuit}}}\]

OpenStudy (anonymous):